FTIR STUDIES OF AMMONIA PHOTOCHEMISTRY IN SOLID PARAHYDROGEN

\underline{MAHMUT RUZI} and DAVID T. ANDERSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

It is believed that producing and trapping high concentrations of reactive species in solid molecular hydrogen will ultimately lead to the development of new high performance rocket fuels.^{a} We think the NH radical could be a viable candidate and try to produce it by photolyzing ammonia (NH_{3}) at low temperature in solid parahydrogen. Upon 193.3 nm photolysis of NH_{3}, we observe both NH_{2} and NH radical photoproducts. No significant changes in the NH radical concentration have been detected during a period of 3 hours at 1.8 K, even though the reaction NH + H_{2} → NH_{3} is highly exothermic and can occur by quantum mechanical tunneling even at these low temperatures. In contrast, the NH_{2} radical is only observed in FTIR scans recorded during photolysis and rapidly decays once the 193.3 nm laser is turned off. We will discuss the possible fates of NH_{2} and how conditions can be optimized to produce high concentrations of NH radicals.

aM. E. Fajardo, S. Tam, T. L. Thompson, and M. E. Cordonnier, Chem. Phys. 189, 351-365 (1994).